# Bridging Gaps in Natural Language Processing for Yorùbá: A Systematic Review of a Decade of Progress and Prospects

### v1

@@ 6.1 Conclusion
This SLR explores NLP progress involving Yorùbá language. It involves surveying studies between $2014$ and $2024$, which have used Yorùbá language in their NLP research, and those with great emphasis on the language, in case of a multilingual setting. Moreover, data were synthesized from these primary studies to deduce findings, thereby procuring answers to the established four research questions, which form the core objectives of the research.
@@
Ultimately, with this study, language researchers will be abreast of the current progress in Yorùbá NLP, thereby equipping them with the necessary ideas to preserve the language through NLP tool representation—this is crucial as it is a widely spoken language with abundant cultural richness and linguistic features. Similarly, it is required to guide future researchers plying this route to eliminate or limit possible odyssey in their research journey.
@@ 6.2 Future Directions
Even though the findings show promising results for NLP research involving Yorùbá language, it is pertinent to outline the current absence of significant efforts in some domains, which are equally important towards improving NLP in the language. For instance, research efforts involving the identification of abusive, offensive, hate speech or cyberbullying, which are all regarded as harmful language usage, have not been explored. This could be due to a lack of or limited annotated corpus in this domain. Such research tasks are essential for ameliorating the usage rate of sensitive words, phrases, and sentences on social media, as they could potentially endanger other users. Consequently, future research efforts can be directed toward building relevant corpora to address these NLP challenges and creating relevant benchmarks to facilitate empirical analysis and continuous research.
Sarcasm detection is another fascinating realm that remains untapped yet for Yorùbá and most under-resource languages. Sarcasm is used to mask the true emotion in a state, mostly by exuding positivity or a seeming positive demeanour. This phenomenon of saying what is not meant or meaning what is not said by natural language users poses challenges for NLP. However, it is essential to accurately detect users’ real emotions for different purposes, including product reviews, feedback analysis, politics & governance, among others. Consequently, developing relevant sarcasm corpora and benchmarks that will potentially birth advanced sarcasm detection models is crucial. Future research can be focused on this domain to promote certainty in natural language usage.
@@
Finally, addressing the drawbacks associated with cultural and social challenges will require developing context-aware models that can adapt to real-world changes. Continuous language usage would also benefit from exploring new cases, such as LLM conversational agents, and developing healthcare and educational tools incorporating NLP. These innovations can aid in bridging the digital gaps among the Yorùbá-speaking communities.
@@ Acknowledgements
No data was used for this research task. However, datasheets that record the different stages in the SLR can be found in the GitHub repository [^17] https://github.com/toheebadura/SLR for reference.

### v2

@@ 6.1 Conclusion
This SLR explores NLP progress involving Yorùbá language. It involves surveying studies between $2014$ and $2024$, which have used Yorùbá language in their NLP research, and those with great emphasis on the language, in case of a multilingual setting. Data were synthesized from these primary studies to deduce findings, thereby providing answers to the established four research questions, forming the research’s core objectives.
@@
In conclusion, with this study, language researchers will be abreast of the current progress in Yorùbá NLP, thereby equipping them with the necessary ideas to preserve the language through NLP tool representation—this is crucial as it is a widely spoken language with abundant cultural richness and linguistic features. Similarly, it is required to guide researchers taking this route in the future to eliminate or limit possible odysseys in their research journey.
@@ 6.2 Future Directions
Even though the findings show promising results for NLP research involving Yorùbá language, it is pertinent to outline the current absence of significant efforts in some domains, which are equally important towards improving NLP in the language. For instance, research efforts involving the identification of abusive, offensive, hate speech, or cyberbullying, which are all regarded as harmful language usage, have limited research as of the time of this publication. This could be due to a lack of or limited annotated corpus in this domain. Such research tasks are essential for ameliorating the usage rate of sensitive words, phrases, and sentences on social media, as they could potentially endanger other users. Consequently, future research efforts can be directed toward building relevant corpora to address these NLP challenges and creating relevant benchmarks to facilitate empirical analysis and continuous research.
Sarcasm detection is another that remains untapped yet for Yorùbá and most under-resourced languages, as noted in the critical analysis compared with the state-of-the-art. Sarcasm is used to mask the true emotion in a state, mostly by exuding positivity or a seeming positive demeanour. This phenomenon of saying what is not meant or meaning what is not said by natural language users poses challenges for NLP. However, it is essential to accurately detect users’ real emotions for different purposes, including product reviews, feedback analysis, politics & governance, among others. Consequently, developing relevant sarcasm corpora and benchmarks that will potentially birth advanced sarcasm detection models is crucial. Future research can be focused on this domain to promote certainty in natural language usage.
@@
Finally, addressing the drawbacks associated with cultural and social challenges requires developing context-aware models that can adapt to real-world changes. Continuous language usage would also benefit from exploring new cases, such as LLM conversational agents, and developing healthcare and educational tools incorporating NLP. These innovations can aid in bridging the digital gaps among the Yorùbá-speaking communities.
@@ Acknowledgements
No data was used for this research task. However, datasheets that record the different stages in the SLR can be found in the GitHub repository [^17] https://github.com/toheebadura/SLR for reference. Also, the published paper can be found online on ScienceDirect [^18] https://doi.org/10.1016/j.nlp.2025.100194.